A pound of alloy of lead and nickel weighs 14.40 ounces in water, where lead losses 1/11 of its weight and nickel losses 1/9 of its weight. How much of each metal is in the alloy?
Answer by Boreal(15235)

x=weight of lead in ounces in alloy out of water, 10/11 x = weight of lead in water
16-x=weight of nickel in ounces in alloy out of water, (8/9)(16-x)=weight of nickel in water
(10x/11)+ (8/9)(16-x)=14.40
(10x/11) + (128/9)-(8x/9)=14.40 multiply everything by 99
90x + 1408-88x=1425.6
2x=17.6
x=8.8 ounces lead
16-x=7.2 ounces nickel
in water the 8.8 ounces of lead weighs 8 ounces
the 7.2 ounces of nickel weighs 6.4 ounces
check